OpenStudy (anonymous):

Three forces of 5n, 7n, and 8n are applied to an object. If the object is in a state of equilibrium: calculate the angle between the lines of action of the 5 N and 7 N forces.

OpenStudy (anonymous):

@agent0smith please help

OpenStudy (agent0smith):

I'll put the 7N force on the x axis just to make it a bit easier.  All the horizontal and vertical components must add up to zero.

OpenStudy (anonymous):

thanks

OpenStudy (agent0smith):

Do you need more help? It looks like there might be a fair bit of work involved...

OpenStudy (agent0smith):

 equating horizontal components gives: \[\Large 8\cos A + 5\cos B = 7\] equating vertical gives: \[\Large 8\sin A = 5\sin B\]

OpenStudy (agent0smith):

Well, that looks like it'll be fun to solve... fair bit of trig. Then you'll find the angle required by 180-B.

OpenStudy (agent0smith):

@eLg were you able to solve it?